Let X be a r.v with pdf f(x) = 3x2 , 0 < x < 1. Then

(a) Calculate the quantities E(X), E(X2) and Var(X).

(b) If the r.v.Y is denoted by Y = 3X − 2, calculate the E(Y ) and Var(Y )


1
Expert's answer
2021-09-16T00:49:15-0400

a)


"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{1}x(3x^2)dx"

"=\big[\dfrac{3x^4}{4}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{3}{4}"



"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{1}x^2(3x^2)dx"

"=\big[\dfrac{3x^5}{5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{3}{5}"

"Var(X)=E(X^2)-(E(X))^2"

"=\dfrac{3}{5}-(\dfrac{3}{4})^2=\dfrac{3}{80}=0.0375"

(b)


"E(Y)=E(3X-2)=3E(X)-2"

"=3(\dfrac{3}{5})-2=-\dfrac{1}{5}=-0.2"

"Var(Y)=Var(3X-2)=(3)^2Var(X)"

"=9(\dfrac{3}{80})=\dfrac{27}{80}=0.3375"